Please help me

Q1. Sections 1 and 2 are at the beginning and end of the bend of the 200 mm diameter pipe in which the quantity of flow is 0.28 m3/s. The angle of deflection of the water is 40. Calculate the force that the liquid exerts on the bend if the pressure in the pipe is 50 kPa. Assume no loss of pressure round the bend.

Q2. A 50 mm diameter stream of water strikes a 0.5 m by 0.5 m square plate which is at an angle of 30 degrees with the stream’s direction. The velocity of the water in the stream is 20 m/s and the jet strikes the door at its centre of gravity. If the plate is hinged at A, by neglecting the friction and self-weight of the plate, determine the normal force P applied at the upper edge of the plate to keep the plate in equilibrium. You may assume that the velocities of the jet remain unchanged after split.
1
Expert's answer
2019-04-12T09:41:35-0400

A1.


"F=ma=m\cdot \Delta v/\Delta t,"

"m=\rho V=\rho Q\Delta t,"

"\Delta v=v_2-v_1=v_1\text{cos}40^\circ-v_1=v_1(\text{cos}40^\circ-1),"




"v_1=Q/A=4Q/(\pi d^2),"


finally,


"F=\rho Q\Delta t\cdot \frac{4Q}{\pi d^2}(\text{cos}40^\circ-1)\frac{1}{\Delta t}="

"=\frac{4\rho Q^2(\text{cos}40^\circ-1)}{\pi d^2}=582\text{ N}."

A2. This problem can be solved likewise. First, calculate the force acting on the centre of the door. To do that, transform the expression for "F" we obtained above considering that "Q=v_1\pi d^2/4":



"F=\frac{4\rho (v_1\pi d^2/4)^2(\text{cos}60^\circ-1)}{\pi d^2}="

"=\frac{\rho v_1^2\pi d^2(\text{cos}60^\circ-1)}{4}=391.9\text{ N}."

Now simply apply equations for moment of force: our force "F" calculated above acts on a centre of plate 50 cm in length:



Assume that the moments are considered in relation to the lower edge.


"0=0.25F-0.5P\text{sin}30^\circ,"

"P=\frac{0.25F}{0.5\text{sin}30^\circ}=391.9\text{ N}."